# InternLM-XComposer2.5-Reward: A Simple Yet Effective Multi-Modal Reward Model

### v1

@@ Abstract
Despite the promising performance of Large Vision Language Models (LVLMs) in visual understanding, they occasionally generate incorrect outputs. While reward models (RMs) with reinforcement learning or test-time scaling offer the potential for improving generation quality, a critical gap remains: publicly available multi-modal RMs for LVLMs are scarce, and the implementation details of proprietary models are often unclear. We bridge this gap with InternLM-XComposer2.5-Reward (IXC-2.5-Reward), a simple yet effective multi-modal reward model that aligns LVLMs with human preferences. To ensure the robustness and versatility of IXC-2.5-Reward, we set up a high-quality multi-modal preference corpus spanning text, image, and video inputs across diverse domains, such as instruction following, general understanding, text-rich documents, mathematical reasoning, and video understanding. IXC-2.5-Reward achieves excellent results on the latest multi-modal reward model benchmark and shows competitive performance on text-only reward model benchmarks. We further demonstrate three key applications of IXC-2.5-Reward: (1) Providing a supervisory signal for RL training. We integrate IXC-2.5-Reward with Proximal Policy Optimization (PPO) yields IXC-2.5-Chat, which shows consistent improvements in instruction following and multi-modal open-ended dialogue; (2) Selecting the best response from candidate responses for test-time scaling; and (3) Filtering outlier or noisy samples from existing image and video instruction tuning training data. To ensure reproducibility and facilitate further research, we have open-sourced all model weights and training recipes at https://github.com/InternLM/InternLM-XComposer.
@@ 1 Introduction
If you don’t know where you are going, you’ll end up some
place else.
@@
Reward Models (RMs) [8, 121, 48, 92, 88, 104, 55, 98, 105, 78, 91] provide the crucial direction guidance about how well an AI model’s outputs align with human preference, and benefit Large Language Models (LLMs) in training and inference. During training, RMs are often used with reinforcement learning from human feedback (RLHF) [66, 7, 71, 70] to penalize undesirable model behaviors and encourage outputs that align with human values. At inference, RMs facilitate test-time scaling strategies [80, 27], such as selecting the best response from candidate outputs or providing step-by-step critiques for complex reasoning tasks [107, 29].
@@
Despite their crucial role in both training and inference, multi-modal RMs for Large Vision Language Models (LVLMs) remain notably underexplored compared to language-only RMs for LLMs. Because current preference data is predominantly text-based and skewed toward specific domains (e.g., safety), data scarcity poses a significant challenge to training multi-modal RMs for diverse modalities such as images, videos, and text. Consequently, existing multi-modal RMs [87, 95] are largely constrained to narrow domains (e.g., mitigating hallucinations) or rely on prompting LVLMs with evaluation prompts, effectively functioning as generative RMs [94]. The limitation of multi-modal RMs subsequently constrains the capabilities of open-source LVLMs such as instruction following and safety-should-refuse, thereby hampering user interaction experience in multi-modal chat scenarios.
@@
The growing community interest in RLHF and test-time scaling highlights the need for multi-modal RMs, which motivates us to present InternLM-XComposer2.5-Reward (IXC-2.5-Reward). Instead of directly transferring unimodal (text) reward models (RMs) to the vision modality, we augment the existing LVLM (InternLM-XComposer2.5) with an additional scoring head to predict reward scores. An effective multi-modal RM should ideally possess two key properties: (1) the ability to predict reward scores for both image, video, and textual inputs and (2) the capacity to generalize across diverse domains, such as instruction following, knowledge, text-rich images (e.g., documents), reasoning tasks, etc. To this end, we develop a pipeline (Fig. 1(a)) to construct multi-modal preference data, and also incorporate existing high-quality datasets. This pipeline selects prompts across diverse domains for text, image, and video inputs, generates corresponding responses, and then uses GPT-4o [31] or verifier [40] to perform preference judgments. Trained on our preference data, IXC-2.5-Reward effectively evaluates both visual (image and video) and textual inputs (Fig. 1 (b)).
@@
IXC-2.5-Reward achieves best performance on multi-modal VL-RewardBench [43] (70.0%) that beat all previous generative RMs including Gemini-1.5-Pro (62.5%) and GPT-4o (62.4%). Even on uni-modal (text) RM benchmarks, IXC-2.5-Reward also demonstrates good results, with an average score of $88.6\%$ on Reward-Bench [41] and $68.8\%$ on RM-Bench [51].
@@
All code and models of IXC-2.5-Reward and IXC-2.5-Chat are publicly available at https://github.com/InternLM/InternLM-XComposer.
Table 1: Overview of existing preference datasets used in IXC-2.5-Reward. IF denotes to Instruction Following.
@@
IF General Tulu-3-IF-augmented-on-policy-8b [40]
UltraFeedback [17]
Safety hhh alignment [5], PKU-Safe [18]
SHP [24], Anthropic-hhrlhf [6]
@@
Chat WildVision-Battle [62]
General LLaVA-Critic [94], VL-Feedback [44],
RLAIF-V [101] MIA-DPO [54]
@@
IF General in-house (will release)
KVQA [76], A-OKVQA [73], PMC-VQA [114]
Text-Rich AI2D [37], IconQA [56], TQA [38]
ChartQA [63], DVQA [36], ScienceQA [57]
Reasoning GeoQA [11], CLEVR-Math [47]
Super-CLEVR [45], TabMWP [58]
@@
General TrafficQA [96], FunQA [93], MiraData [35]
@@ 2 Related Work
Reward models (RMs) are crucial for both Reinforcement Learning from Human Feedback (RLHF) [66, 7] and Test-time Scaling Laws [80, 29]. RMs have different implementation forms, such as (1) discriminative RM [8, 121, 48, 92, 88, 104, 55, 98], usually a sequence classifier that classifies input sequences into different categories, such as binary classification (“good” or “bad,”) or on a more granular scale [92, 88]. (2) generative RM [39, 105, 78, 91] that are prompted to generate the feedback in the form of text, often a critique or explanation of why a certain output is good or bad. (3) implicit RMs [32, 40] that are models optimized using DPO [70] that the predicted log probabilities are interpreted as implicit reward signal. Besides, RMs can also be divided into Outcome RMs (ORMs) [16] and Process RMs (PRMs) [86, 46, 75]. Our IXC-2.5-Reward belongs to the discriminative RM and ORM.
Reward Model in Large Vision-Language Models
@@
Previous RMs for LVLMs [87, 94, 95] are limited to specific domains (e.g., reducing hallucination) or developed using relatively weak base models, which makes the implemented models significantly inferior to LLM RMs. The lack of effective multi-modal RMs has created a bottleneck in vision RLHF, forcing researchers to merely use the variants of the off-poly DPO algorithm [70]. Previous work using open-source LVLMs as generative RMs [102, 65, 95], injection of hallucinations with data augmentation techniques [22, 25, 120, 122, 67, 33, 21] and rule-based selection [9, 53] for DPO data selection, which potentially compromise performance compared to the on-policy RL solutions like PPO [71]. Moreover, lacking multi-modal RMs has also led to the reliance on human annotation [82, 100] or the use of proprietary models [109, 116] like GPT4 as generative RMs for DPO pair selection, which is expensive and unsustainable for large-scale applications. Although open-source RMs for LVLMs have lagged behind their LLM counterparts, the growing community interest highlights the need for multi-modal RMs, which motivates our work. In this work, we demonstrate that IXC-2.5-Reward is capable of combining with the PPO training and for DPO data selection at a low cost.
Reward Model Evaluations
@@
The development of evaluation benchmarks is essential for improving RMs. Several comprehensive benchmarks have been proposed for evaluating RMs of LLMs, such as general abilities [41, 119, 51], multilingual [81, 28], RAG [34], and mathematical process reward [117]. The limited availability of multi-modal RMs has hampered the development of evaluation benchmarks, with existing benchmark [43] focusing solely on generative RMs and lacking the evaluation of process supervision. However, given the critical importance of RMs, we expect significant progress in multi-modal RM benchmarking in the future.
@@
Data Preparation
Reward models are trained using pairwise preference annotations (e.g., prompts $x$ with chosen responses $y_{c}$ and rejected responses $y_{r}$) that reflect human preferences. While existing public preference data is primarily textual, with limited image and scarce video examples, we train IXC-2.5-Reward using both open-source data and a newly collected dataset to ensure broader domain coverage.
Tab. 2 lists the open-source pairwise data used in IXC-2.5-Reward, primarily focused on instruction following, safety, and general knowledge. Tab. 2 details the source of our newly collected data, which is initially the supervised fine-tuning (SFT) data consisting of prompts $x$ and corresponding chosen responses $y_{c}$ across diverse domains: text-rich document understanding, math reasoning, and video understanding. We also collect some in-house data about the instruction following, which will be released in the future. To obtain rejected responses $y_{r}$, we prompt the SFT model, InternLM-XComposer-2.5 (IXC-2.5) [112] to generate multiple outputs for each prompt and then employ distinct selection criteria. For general and text-rich data, we use GPT-4o [31] with pairwise evaluation prompts to determine the rejected response that was evaluated worse than the SFT ground-truth answer. For math reasoning and instruction following data, we build verifier functions [40] that compare generated responses against ground-truth solutions to label the chosen and rejected data. Our newly collected data complements existing open-source data, creating a comprehensive, high-quality multi-modal preference dataset.
Model Architecture
Our reward model InternLM-XComposer 2.5-Reward (IXC-2.5-Reward) is built upon the SFT model (IXC-2.5) [111]. As shown in Fig. 1 (b), we use the pre-trained weights of IXC-2.5-Chat for most of the parts, such as the visual encoder and the MLP projector, which has aligned the image and video data with text modalities. Thus, the IXC-2.5-Reward is merely required to train preference data to predict the reward score and avoid using other pre-training data for modality alignment.
@@
Loss Function
Our reward model is trained via the following loss function:
$\mathcal{L}_{\text{RM}}=-E(\log(\sigma(r(x,y_{w}))-r(x,y_{l}))),$ (1)
@@
Training Strategy
As shown in Fig. 1 (b), we froze the model’s vision encoder and projector that are initialized from IXC-2.5 [112], training only the LLM (InternLM [112]) and the score head. Other components of IXC-2.5, such as the dynamic image partitioning mechanism for high-resolution inputs, remained unchanged.
Length Constraints
We remove data pairs where the length of the chosen response $y_{w}$ is significantly longer than the length of the rejected response $y_{l}$. This helps prevent the reward model from learning to associate length with quality. Notably, we found that the vulnerability of LLM-based evaluation to length bias, a known issue in LLMs [23], has also significant implications for LVLMs. Specifically, open-ended Visual Question Answering (VQA) benchmarks that employ LVLMs (e.g., GPT-4o) as judges are susceptible to inflated scores from overly long responses. Consequently, removing the length constraint on the reward model resulted in improved PPO policy performance. A detailed analysis is provided in Tab. 7.
Figure 2: Using IXC-2.5-Reward for Data Cleaning. We visualize the outlier and noisy examples detected by IXC-2.5-Reward with low reward scores from existing image and video instruction-tuning datasets, such as ALLaVA [10] and LLaVA-Video-178K [115]. The “Explain” refers to explanations of error causes as identified by human experts, rather than outputs generated by the reward model.
Table 3: Evaluation results on VLRewardBench [43]. The best and second-best results for proprietary models and open-source models are highlighted in bold and underlined, respectively.
@@
Gemini-1.5-Flash (2024-09-24) [83] - 47.8 59.6 58.4 57.6 55.3
Gemini-1.5-Pro (2024-09-24) [83] - 50.8 72.5 64.2 67.2 62.5
Claude-3.5-Sonnet (2024-06-22) [4] - 43.4 55.0 62.3 55.3 53.6
GPT-4o-mini (2024-07-18) [3] - 41.7 34.5 58.2 41.5 44.8
GPT-4o (2024-08-06) [3] - 49.1 67.6 70.5 65.8 62.4
@@
LLaVA-OneVision-7B-ov [42] 7B 32.2 20.1 57.1 29.6 36.5
Qwen2-VL-7B [90] 7B 31.6 19.1 51.1 28.3 33.9
Molmo-7B [20] 7B 31.1 31.8 56.2 37.5 39.7
InternVL2-8B [85] 8B 35.6 41.1 59.0 44.5 45.2
LLaVA-Critic-8B [94] 8B 54.6 38.3 59.1 41.2 44.0
Llama-3.2-11B [84] 11B 33.3 38.4 56.6 42.9 42.8
Pixtral-12B [1] 12B 35.6 25.9 59.9 35.8 40.4
Molmo-72B [20] 72B 33.9 42.3 54.9 44.1 43.7
Qwen2-VL-72B [90] 72B 38.1 32.8 58.0 39.5 43.0
NVLM-D-72B [19] 72B 38.9 31.6 62.0 40.1 44.1
Llama-3.2-90B [84] 90B 42.6 57.3 61.7 56.2 53.9
@@
Having the reward model IXC-2.5-Reward enables the application of on-policy reinforcement learning algorithms (e.g., PPO [71], RLOO [2], GRPO [77]) to optimize LVLM performance towards desired human preferences directly. Using the PPO [71] algorithm, we subsequently train the policy model (IXC-2.5-Chat, $\pi_{\theta}$) to maximize expected rewards from our reward model (IXC-2.5-Reward) while staying close to the reference model (IXC-2.5, $\pi_{\text{ref}}$) for stability. A critic model $V$, initialized from IXC-2.5-Reward, is trained alongside $\pi_{\theta}$ to reduce the variance of policy updates.
Data Prepration
Similar to findings in [30], we found that average reward scores differ across task domains (e.g., general, text-rich, reasoning). This work focuses on improving the policy model’s instruction following and open-ended chat abilities, which are critical for real-world applications such as stream chatting and human-AI interaction [110]. Simultaneously, we ensure that performance in other domains (e.g., text-rich, reasoning) is not degraded relative to the SFT model IXC-2.5. Using our multi-modal preference data (which trains IXC-2.5-Reward), we curate a prompt set that prioritizes general chat and instruction following, while ensuring diversity through the inclusion of text-rich documents, math reasoning, and video understanding.
PPO
The PPO training begins by sampling a prompt from our prompt set. Then, the policy $\theta_{\pi}$ model generates responses, and the reward model computes the reward score $r_{t}$ at each state $s_{t}$ at the time-step $t$. Given the reward score $r_{t}$ and and the critic model $V$, we compute the temporal difference error $\delta_{t}$, the Generalized Advantage Estimation (GAE) [72] $A_{t}$, and the Returns $R_{t}$ as:
@@
Garbage in, garbage out. Problematic samples in instruction tuning datasets negatively impact LVLM training. While existing methods [13] employ classifiers like CLIP [69] for filtering, these approaches have limitations, particularly with long-context inputs [108], high-resolution images, or videos. As shown in Fig. 2, we observe a strong correlation between low IXC-2.5-Reward scores and problematic samples, including hallucinations, empty answers, and irrelevant image/video-text pairings. Therefore, IXC-2.5-Reward effectively cleans both pre-training and post-training data for LVLMs.

### v2

@@ Abstract
Despite the promising performance of Large Vision Language Models (LVLMs) in visual understanding, they occasionally generate incorrect outputs. While reward models (RMs) with reinforcement learning or test-time scaling offer the potential for improving generation quality, a critical gap remains: publicly available multi-modal RMs for LVLMs are scarce, and the implementation details of proprietary models are often unclear. We bridge this gap with InternLM-XComposer2.5-Reward (IXC-2.5-Reward), a simple yet effective multi-modal reward model that aligns LVLMs with human preferences. To ensure the robustness and versatility of IXC-2.5-Reward, we set up a high-quality multi-modal preference corpus spanning text, image, and video inputs across diverse domains, such as instruction following, general understanding, text-rich documents, mathematical reasoning, and video understanding. IXC-2.5-Reward achieves excellent results on the latest multi-modal reward model benchmark and shows competitive performance on text-only reward model benchmarks. We further demonstrate three key applications of IXC-2.5-Reward: (1) Providing a supervisory signal for RL training. We integrate IXC-2.5-Reward with Proximal Policy Optimization (PPO) yields IXC-2.5-Chat, which shows consistent improvements in instruction following and multi-modal open-ended dialogue; (2) Selecting the best response from candidate responses for test-time scaling; and (3) Filtering outlier or noisy samples from existing image and video instruction tuning training data.
InternLM-XComposer2.5-Reward: A Simple Yet Effective
Multi-Modal Reward Model
@@ 1 Introduction
If you don’t know where you are going, you’ll end
up some place else.
@@
Reward Models (RMs) Cai et al. (2024); Zhu et al. (2023); Liu et al. (2024a); Wang et al. (2024f, b); Yuan et al. (2024a); Lou et al. (2024); Yang et al. (2024b); Yuan et al. (2024b); Shiwen et al. (2024); Wang et al. (2024e) provide the crucial direction guidance about how well an AI model’s outputs align with human preference, and benefit Large Language Models (LLMs) in training and inference. During training, RMs are often used with reinforcement learning from human feedback (RLHF) Ouyang et al. (2022); Bai et al. (2022b); Schulman et al. (2017); Rafailov et al. (2024) to penalize undesirable model behaviors and encourage outputs that align with human values. At inference, RMs facilitate test-time scaling strategies Snell et al. (2024); Gulcehre et al. (2023), such as selecting the best response from candidate outputs or providing step-by-step critiques for complex reasoning tasks Zelikman et al. (2022); Hosseini et al. (2024).
@@
Despite their crucial role in both training and inference, multi-modal RMs for Large Vision Language Models (LVLMs) remain notably underexplored compared to language-only RMs for LLMs. Because current preference data is predominantly text-based and skewed toward specific domains (e.g., safety), data scarcity poses a significant challenge to training multi-modal RMs for diverse modalities such as images, videos, and text. Consequently, existing multi-modal RMs Wang et al. (2024a); Xiyao et al. (2024) are largely constrained to narrow domains (e.g., mitigating hallucinations) or rely on prompting LVLMs with evaluation prompts, effectively functioning as generative RMs Xiong et al. (2024). The limitation of multi-modal RMs subsequently constrains the capabilities of open-source LVLMs such as instruction following and safety-should-refuse, thereby hampering user interaction experience in multi-modal chat scenarios.
@@
The growing community interest in RLHF and test-time scaling highlights the need for multi-modal RMs, which motivates us to present InternLM-XComposer2.5-Reward (IXC-2.5-Reward). Instead of directly transferring unimodal (text) reward models (RMs) to the vision modality, we augment the existing LVLM (InternLM-XComposer2.5) with an additional scoring head to predict reward scores. An effective multi-modal RM should ideally possess two key properties: (1) the ability to predict reward scores for both image, video, and textual inputs and (2) the capacity to generalize across diverse domains, such as instruction following, knowledge, text-rich images (e.g., documents), reasoning tasks, etc. To this end, we develop a pipeline (Fig. 1(a)) to construct multi-modal preference data, and also incorporate existing high-quality datasets. This pipeline selects prompts across diverse domains for text, image, and video inputs, generates corresponding responses, and then uses GPT-4o Hurst et al. (2024) or verifier Lambert et al. (2024a) to perform preference judgments. Trained on our preference data, IXC-2.5-Reward effectively evaluates both visual (image and video) and textual inputs (Fig. 1 (b)).
@@
IXC-2.5-Reward achieves best performance on multi-modal VL-RewardBench Li et al. (2024b) (70.0%) that beat all previous generative RMs including Gemini-1.5-Pro (62.5%) and GPT-4o (62.4%). Even on uni-modal (text) RM benchmarks, IXC-2.5-Reward also demonstrates good results, with an average score of $88.6\%$ on Reward-Bench Lambert et al. (2024b) and $68.8\%$ on RM-Bench Liu et al. (2024b).
@@
Table 1: Overview of existing preference datasets used in IXC-2.5-Reward.
@@
IF General Tulu-3-IF-augmented-on-policy-8b Lambert et al. (2024a)
UltraFeedback Cui et al. (2024)
Safety hhh alignment Askell et al. (2021), PKU-Safe Dai et al. (2024a)
SHP Ethayarajh et al. (2022), Anthropic-hhrlhf Bai et al. (2022a)
@@
Chat WildVision-Battle Lu et al. (2024c)
General LLaVA-Critic Xiong et al. (2024), VL-Feedback Li et al. (2024c),
RLAIF-V Yu et al. (2024b) MIA-DPO Liu et al. (2024e)
@@
IF General MM-IFDPO-23k Ding et al. (2025)
KVQA Shah et al. (2019), A-OKVQA Schwenk et al. (2022), PMC-VQA Zhang et al. (2023)
Text-Rich AI2D Kembhavi et al. (2016), IconQA Lu et al. (2021), TQA Kembhavi et al. (2017)
ChartQA Masry et al. (2022), DVQA Kafle et al. (2018), ScienceQA Lu et al. (2022a)
Reasoning GeoQA Chen et al. (2021), CLEVR-Math Lindström and Abraham (2022)
Super-CLEVR Li et al. (2023), TabMWP Lu et al. (2022b)
@@
General TrafficQA Xu et al. (2021), FunQA Xie et al. (2024), MiraData Ju et al. (2024)
@@ 2 Related Work
Reward Model in Large Language Models. Reward models (RMs) are crucial for both Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022); Bai et al. (2022b) and Test-time Scaling Laws Snell et al. (2024); Hosseini et al. (2024). RMs have different implementation forms, such as (1) discriminative RM Cai et al. (2024); Zhu et al. (2023); Liu et al. (2024a); Wang et al. (2024f, b); Yuan et al. (2024a); Lou et al. (2024); Yang et al. (2024b), usually a sequence classifier that classifies input sequences into different categories, such as binary classification (“good” or “bad,”) or on a more granular scale Wang et al. (2024f, b). (2) generative RM Kim et al. (2023); Yuan et al. (2024b); Shiwen et al. (2024); Wang et al. (2024e) that are prompted to generate the feedback in the form of text, often a critique or explanation of why a certain output is good or bad. (3) implicit RMs Ivison et al. (2023); Lambert et al. (2024a) that are models optimized using DPO Rafailov et al. (2024) that the predicted log probabilities are interpreted as implicit reward signal. Besides, RMs can also be divided into Outcome RMs (ORMs) Cobbe et al. (2021) and Process RMs (PRMs) Uesato et al. (2022); Lightman et al. (2023); Setlur et al. (2024). Our IXC-2.5-Reward belongs to the discriminative RM and ORM.
@@
Reward Model in Large Vision-Language Models. Previous RMs for LVLMs Wang et al. (2024a); Xiong et al. (2024); Xiyao et al. (2024) are limited to specific domains (e.g., reducing hallucination) or developed using relatively weak base models, which makes the implemented models significantly inferior to LLM RMs. The lack of effective multi-modal RMs has created a bottleneck in vision RLHF, forcing researchers to merely use the variants of the off-poly DPO algorithm Rafailov et al. (2024). Previous work using open-source LVLMs as generative RMs Yu et al. (2024c); Ouali et al. (2025); Xiyao et al. (2024), injection of hallucinations with data augmentation techniques Deng et al. (2024b); Favero et al. (2024); Zhou et al. (2024b); Zhu et al. (2024); Pi et al. (2025); Jiang et al. (2024); Deng et al. (2024a) and rule-based selection Cao et al. (2024); Liu et al. (2024f) for DPO data selection, which potentially compromise performance compared to the on-policy RL solutions like PPO Schulman et al. (2017). Moreover, lacking multi-modal RMs has also led to the reliance on human annotation Sun et al. (2023); Yu et al. (2024a) or the use of proprietary models Zhang et al. (2024a); Zhao et al. (2023) like GPT4 as generative RMs for DPO pair selection, which is expensive and unsustainable for large-scale applications. Although open-source RMs for LVLMs have lagged behind their LLM counterparts, the growing community interest highlights the need for multi-modal RMs, which motivates our work. In this work, we demonstrate that IXC-2.5-Reward is capable of combining with the PPO training and for DPO data selection at a low cost.
@@
Reward Model Evaluations. The development of evaluation benchmarks is essential for improving RMs. Several comprehensive benchmarks have been proposed for evaluating RMs of LLMs, such as general abilities Lambert et al. (2024b); Zhou et al. (2024a); Liu et al. (2024b), multilingual Son et al. (2024); Gureja et al. (2024), RAG Jin et al. (2024), and mathematical process reward Zheng et al. (2024). The limited availability of multi-modal RMs has hampered the development of evaluation benchmarks, with existing benchmark Li et al. (2024b) focusing solely on generative RMs and lacking the evaluation of process supervision. However, given the critical importance of RMs, we expect significant progress in multi-modal RM benchmarking in the future.
@@
Data Preparation. Reward models are trained using pairwise preference annotations (e.g., prompts $x$ with chosen responses $y_{c}$ and rejected responses $y_{r}$) that reflect human preferences. While existing public preference data is primarily textual, with limited image and scarce video examples, we train IXC-2.5-Reward using both open-source data and a newly collected dataset to ensure broader domain coverage.
Tab. 2 lists the open-source pairwise data used in IXC-2.5-Reward, primarily focused on instruction following, safety, and general knowledge. Tab. 2 details the source of our newly collected data, which is initially the supervised fine-tuning (SFT) data consisting of prompts $x$ and corresponding chosen responses $y_{c}$ across diverse domains: text-rich document understanding, math reasoning, and video understanding. We also collect some in-house data about the instruction following, which will be released in the future. To obtain rejected responses $y_{r}$, we prompt the SFT model, InternLM-XComposer-2.5 (IXC-2.5) Zhang et al. (2024c) to generate multiple outputs for each prompt and then employ distinct selection criteria. For general and text-rich data, we use GPT-4o Hurst et al. (2024) with pairwise evaluation prompts to determine the rejected response that was evaluated worse than the SFT ground-truth answer. For math reasoning and instruction following data, we build verifier functions Lambert et al. (2024a) that compare generated responses against ground-truth solutions to label the chosen and rejected data. Our newly collected data complements existing open-source data, creating a comprehensive, high-quality multi-modal preference dataset.
Model Architecture. Our reward model InternLM-XComposer 2.5-Reward (IXC-2.5-Reward) is built upon the SFT model (IXC-2.5) Zhang et al. (2024d). As shown in Fig. 1 (b), we use the pre-trained weights of IXC-2.5-Chat for most of the parts, such as the visual encoder and the MLP projector, which has aligned the image and video data with text modalities. Thus, the IXC-2.5-Reward is merely required to train preference data to predict the reward score and avoid using other pre-training data for modality alignment.
@@
Loss Function. Our reward model is trained via the following loss function:
$\mathcal{L}_{\text{RM}}=-E(\log(\sigma(r(x,y_{w})-r(x,y_{l})))),$ (1)
@@
Training Strategy. As shown in Fig. 1 (b), we froze the model’s vision encoder and projector that are initialized from IXC-2.5 Zhang et al. (2024c), training only the LLM (InternLM Zhang et al. (2024c)) and the score head. Other components of IXC-2.5, such as the dynamic image partitioning mechanism for high-resolution inputs, remained unchanged.
Length Constraints. We remove data pairs where the length of the chosen response $y_{w}$ is significantly longer than the length of the rejected response $y_{l}$. This helps prevent the reward model from learning to associate length with quality. Notably, we found that the vulnerability of LLM-based evaluation to length bias, a known issue in LLMs Dubois et al. (2024), has also significant implications for LVLMs. Specifically, open-ended Visual Question Answering (VQA) benchmarks that employ LVLMs (e.g., GPT-4o) as judges are susceptible to inflated scores from overly long responses. Consequently, removing the length constraint on the reward model resulted in improved PPO policy performance. A detailed analysis is provided in Tab. 7.
Figure 2: Using IXC-2.5-Reward for Data Cleaning. We visualize the outlier and noisy examples detected by IXC-2.5-Reward with low reward scores from existing image and video instruction-tuning datasets, such as ALLaVA Chen et al. (2024a) and LLaVA-Video-178K Zhang et al. (2024e). The “Explain” refers to explanations of error causes as identified by human experts, rather than outputs generated by the reward model.
Table 3: Evaluation results on VLRewardBench Li et al. (2024b). The best and second-best results for proprietary models and open-source models are highlighted in bold and underlined, respectively.
@@
Gemini-1.5-Flash (2024-09-24) Team (2024a) - 47.8 59.6 58.4 57.6 55.3
Gemini-1.5-Pro (2024-09-24) Team (2024a) - 50.8 72.5 64.2 67.2 62.5
Claude-3.5-Sonnet (2024-06-22) Anthropic (2024) - 43.4 55.0 62.3 55.3 53.6
GPT-4o-mini (2024-07-18) AI (2024) - 41.7 34.5 58.2 41.5 44.8
GPT-4o (2024-08-06) AI (2024) - 49.1 67.6 70.5 65.8 62.4
@@
LLaVA-OneVision-7B-ov Li et al. (2024a) 7B 32.2 20.1 57.1 29.6 36.5
Qwen2-VL-7B Wang et al. (2024d) 7B 31.6 19.1 51.1 28.3 33.9
Molmo-7B Deitke et al. (2024) 7B 31.1 31.8 56.2 37.5 39.7
InternVL2-8B Team (2024c) 8B 35.6 41.1 59.0 44.5 45.2
LLaVA-Critic-8B Xiong et al. (2024) 8B 54.6 38.3 59.1 41.2 44.0
Llama-3.2-11B Team (2024b) 11B 33.3 38.4 56.6 42.9 42.8
Pixtral-12B Agrawal et al. (2024) 12B 35.6 25.9 59.9 35.8 40.4
Molmo-72B Deitke et al. (2024) 72B 33.9 42.3 54.9 44.1 43.7
Qwen2-VL-72B Wang et al. (2024d) 72B 38.1 32.8 58.0 39.5 43.0
NVLM-D-72B Dai et al. (2024b) 72B 38.9 31.6 62.0 40.1 44.1
Llama-3.2-90B Team (2024b) 90B 42.6 57.3 61.7 56.2 53.9
@@
Having the reward model IXC-2.5-Reward enables the application of on-policy reinforcement learning algorithms (e.g., PPO Schulman et al. (2017), RLOO Ahmadian et al. (2024), GRPO Shao et al. (2024)) to optimize LVLM performance towards desired human preferences directly. Using the PPO Schulman et al. (2017) algorithm, we subsequently train the policy model (IXC-2.5-Chat, $\pi_{\theta}$) to maximize expected rewards from our reward model (IXC-2.5-Reward) while staying close to the reference model (IXC-2.5, $\pi_{\text{ref}}$) for stability. A critic model $V$, initialized from IXC-2.5-Reward, is trained alongside $\pi_{\theta}$ to reduce the variance of policy updates.
Data Prepration. Similar to findings in Hou et al. (2024), we found that average reward scores differ across task domains (e.g., general, text-rich, reasoning). This work focuses on improving the policy model’s instruction following and open-ended chat abilities, which are critical for real-world applications such as stream chatting and human-AI interaction Zhang et al. (2024b). Simultaneously, we ensure that performance in other domains (e.g., text-rich, reasoning) is not degraded relative to the SFT model IXC-2.5. Using our multi-modal preference data (which trains IXC-2.5-Reward), we curate a prompt set that prioritizes general chat and instruction following, while ensuring diversity through the inclusion of text-rich documents, math reasoning, and video understanding.
PPO. The PPO training begins by sampling a prompt from our prompt set. Then, the policy $\theta_{\pi}$ model generates responses, and the reward model computes the reward score $r_{t}$ at each state $s_{t}$ at the time-step $t$. Given the reward score $r_{t}$ and and the critic model $V$, we compute the temporal difference error $\delta_{t}$, the Generalized Advantage Estimation (GAE) Schulman et al. (2018) $A_{t}$, and the Returns $R_{t}$ as:
@@
Garbage in, garbage out. Problematic samples in instruction tuning datasets negatively impact LVLM training. While existing methods Chen et al. (2024c) employ classifiers like CLIP Radford et al. (2021) for filtering, these approaches have limitations, particularly with long-context inputs Zhang et al. (2025a), high-resolution images, or videos. As shown in Fig. 2, we observe a strong correlation between low IXC-2.5-Reward scores and problematic samples, including hallucinations, empty answers, and irrelevant image/video-text pairings. Therefore, IXC-2.5-Reward effectively cleans both pre-training and post-training data for LVLMs.